OpenStudy (anonymous):

8th grade math, plzz help!! A fruit stand sells melons, pears and grapes in the ratio of 2:3:5, respectively. If the stand sold 45 pears, what is the total number of melons and grapes sold? *Show all work*

OpenStudy (anonymous):

45 / 3 = 15 then 15 * 2 = 30 for melons, and 15 * 5 = 75 for grapes.